solve both please
Sketch the region enclosed by y = 3x and y 6.0². Find the area of the region.
Find the average value of the function f(x) 3x4 on the interval 0 < x < 2
Answer #1

9=30, y=612. thu NOW com paring two equation aboy 31 = 6x2 61-3X=0 a (61-3)=0 1-0 6-1-320 260 613 fa Blon 150 X0 12 1/2 Areaf(1) = 3x4 OS232 Average value or a function fi in alasb b favo s fuldx bra a 2 - J 3x4 de 2-0 0 2 do *3 is 0 2 4+) 2 3 4+1 0